OpenStudy (anonymous):

Lijun Wang and colleagues at Princeton sent a laser pulse through a gas containing cesium atoms. The light beam traveled through the cesium atoms at a speed of 9.3 × 1010 meters per second. Normally, the speed of light is 3.0 × 108 meters per second. How many times faster than normal was the speed of light in the cesium gas?

OpenStudy (anonymous):

\[9.3\times 10^{10}\div 3\times 10^{8}\]

OpenStudy (anonymous):

\[3.3\times 10^{10-8}=3.3\times 10^2=3.3\times 100=330\]

OpenStudy (mertsj):

Psst (3.1)